OpenStudy (anonymous):

log 8 of base to to the 33rd power??? Please help. i know how to begin, but i don't know how to finish it log(2)8^33

OpenStudy (anonymous):

\[\log _{2} 8^{33} \]

OpenStudy (anonymous):

wat is the question?

OpenStudy (anonymous):

evaluate the expression

OpenStudy (anonymous):

wat is the formula for converting any base to base 10 again?

OpenStudy (anonymous):

i know that it will be \[33\log _{2}8\]

OpenStudy (anonymous):

Use the fact that:\[\log_{b}{b}=1\]If you can somehow write \[\log_{2}{8}\]\[\log_{2}{2^n}\]for some power n, you will be done.

OpenStudy (anonymous):

i could solve this if there wasn't 33

OpenStudy (anonymous):

pretend the 33 isnt there then. Can you solve\[\log_{2}{8}\]? That will get you closer to the answer.

OpenStudy (anonymous):

3

OpenStudy (anonymous):

Right. So use that property you used earlier, in combination with the answer you just gave:\[\log_{2}{8^{33}}=33\cdot\left(\log_{2}{8}\right)=33\cdot 3\]

OpenStudy (anonymous):

so, should i only multiply 3x33?

OpenStudy (anonymous):

i got it! Thank you sooooooo much!!!!!!!

OpenStudy (anonymous):

Joe, thanks again!